one complementary angle is 10 degrees more than the second. find the two angles.
Answer by checkley79(3341)

The sum of two complementary angles equal 90 degrees.
x+(x+10)=90
2x=90-10
2x=80
x=80/2
x=40 degrees for the smaller angle.
40+10=50 degrees for the larger angle.
Proof:
40+50=90
90=90